# Can Emulating Semantic Translation Help LLMs with Code Translation? A Study Based on Pseudocode

### v2

@@ 3.1. Research Questions
This RQ aims to show a general picture of the effectiveness of semantic translation for code via pseudocode. We also investigate how such pseudocode-based code translation can work with the widely adopted direct code translation, e.g., fully replacing direct translation or complementing it by adopting both strategies simultaneously.
RQ2: How effective is pseudocode-based semantic translation for different pairs of source and target programming languages?
@@
RQ3: Is pseudocode a more effective intermediary compared to a specific programming language?
Macedo et al. (2025) demonstrates that a specific PL with close syntactic and semantic similarities to the source and target PLs may act as an effective transitive intermediary to benefit LLMs in code translation. For example, translating Python code to Java via Rust is found to lead to better accuracy than direct translation. In this RQ, we aim to understand whether pseudocode can serve as a more effective intermediary than a specific transitive PL.
@@
LLMs may not accurately interpret and record the intent of the input program, resulting in low-quality pseudocode that may conversely hinder LLMs from generating correct code translation. This RQ investigates the code translation performance based on high-quality pseudocode, aiming to reveal the potential of pseudocode-based code translation and identify the bottlenecks of studied LLMs to harness the potential effectiveness.
@@ 3.2.1. Programming Languages
In this study, we evaluate code translation performance across six widely studied PLs, i.e., C++, Python, Java, JavaScript, Go, and Rust. The choice of these PLs follows prior work, where translation among the first four PLs has been extensively explored (Rozière et al., 2022; Jiao et al., 2023; Yang et al., 2024), while recent studies have highlighted the emerging PLs such as Go and Rust (Pan et al., 2024; Macedo et al., 2025).
@@ 3.2.2. Code Translation Tasks
We prompt LLMs to translate the solution programs of 323 LeetCode problems in LiveCodeBench (Jain et al., 2025). We choose to translate such programs since they involve common implementations widely used in practical development and become a representative benchmark to evaluate the coding ability of LLMs (Hui et al., 2024; Team, 2024). Actually, solution programs of LeetCode problems have been widely adopted as the subjects in code translation studies since they formulate easy-to-collect multilingual parallel programs (Wang et al., 2023a; Ahmad et al., 2023). Moreover, LiveCodeBench provides carefully validated and representative tests, facilitating the reliable validation of the translation results. Additionally, the LiveCodeBench programming problems are released after May 2023 (Jain et al., 2025), allowing the evaluation to be less subject to the data contamination issue compared to the conventional code translation benchmarks like CodeNet (Puri et al., 2021), TransCoder (Rozière et al., 2020), and AVATAR (Ahmad et al., 2023) established in 2020–2021.
To support the investigation of code translation among diverse PLs, we follow existing practices (Jain et al., 2025; Wu et al., 2025) to collect correct user-written solution programs that can pass all the tests on LiveCodeBench, for all six PLs. These solution programs are used as the source programs to translate. Among all 381 LeetCode problems from LiveCodeBench, we excluded 16 problems with incorrect tests identified by Wu et al. (2025). We also excluded another 42 problems lacking solutions in at least one of the six PLs on LeetCode. This results in 116 easy problems, 160 medium problems, and 47 hard problems. Note that, for each problem, LiveCodeBench does not provide corresponding solution programs and only supports the evaluation of Python solutions. We collect the solution programs in six PLs and extend the evaluation scaffold to all six studied PLs via execution-based validation, following Wu et al. (2025) (who extended the support of C++ and Rust). We open-source these materials to facilitate future research on code translation (Artifact of this paper, [n. d.]).
@@ 3.3. Experimental Strategies for Code Translation
    Direct Translation (abbr. $\mathcal{D}$): This strategy prompts LLMs to translate the programs in the source PL to the target PL. The strategy is widely adopted by existing LLM-based code translation studies (Yang et al., 2024; Pan et al., 2024). We follow these studies to design the prompt as shown in Figure 3(a).
@@
    Transitive Translation via Pseudocode (abbr. $\mathcal{P}$): This strategy prompts LLMs to explicitly conduct semantic translation using two transitive steps: (i) generate a piece of pseudocode based on the original program following the prompt from Wu et al. (2025); and (ii) implement the pseudocode into a translated program in the target PL following the prompt in Figure 3(b). As introduced in Section 2.2, this strategy is inspired by human practices of translation and the findings that pseudocode can encode the PL-agnostic code intent and logic of a program (Wu et al., 2025), serving as an appropriate semantic representation for emulating semantic translation. In our study, pseudocode and translated programs are generated by the same LLM (referred to as “translator LLM”). The prompts are designed following the practices of Wu et al. (2025).
@@
    Transitive Translation via Pseudocode with Original Program as Context (abbr. $\mathcal{PC}$): This strategy extends the above pseudocode-based strategy to include more implementation details (i.e., original program) as the context to guide code translation. Specifically, considering that pseudocode abstracts implementation details of the original program, it may miss some helpful information for LLMs to implement the original program’s intent and logic in the target PL. Thus, we include the original program in the source PL as an implementation example of the pseudocode for LLMs’ reference when generating the target code through explicit semantic translation, using the prompt shown in Figure 3(c).
@@
We also examine the effectiveness of combining direct translation with pseudocode-based translation. Specifically, we design two straightforward hybrid strategies by evenly mixing translation results of direct translation and transitive translation via pseudocode, while keeping the same total attempts as the above strategies (i.e., 10 attempts as mentioned in Section 2.1, with five attempts for each approach). The two hybrid strategies are defined as follows:
@@
(a) Prompt for Direct Translation
(b) Prompt for Generating Code from Pseudocode
(c) Prompt for Generating Code from Pseudocode with Source Program as Reference Implementation
@@ 3.4. Experimental Setups
Metric: We report the pass@10 rate of the translated programs (i.e., programs in the target PL) on the test cases provided by LiveCodeBench to reflect the computational accuracy of the translation results. The metric measures the ratio of tasks on which the translator LLM successfully generates a correct translation in at least one out of ten attempts. Compared with text-based metrics like BLEU and Exact Match rate, the execution-based pass@10 metric can measure the semantic equivalence of the translation results based on the high-quality tests provided in LiveCodeBench. The metric is widely adopted in code generation (Jain et al., 2025; Wu et al., 2025) and translation studies (Macedo et al., 2025). Note that we use pass@10 rather than using pass@1 for one attempt, because we aim to investigate the maximum potential of the studied strategies and LLMs in code translation, considering the higher translation accuracy achieved with ten attempts than one single attempt demonstrated by existing study (Macedo et al., 2025), as well as the practical feasibility of filtering out the incorrect translations based on the source program’s behavior as discussed in Section 2.1. The result of pass@1 shows similar trends.
Studied LLMs: In this study, we investigate five popular LLMs from different famous LLM families deployable on our machine. They are Qwen2.5-Coder-32B-Instruct (abbr. Qwen32B) (Hui et al., 2024), Qwen2.5-Coder-7B-Instruct (abbr. Qwen7B) (Hui et al., 2024), Phi-4 (Abdin et al., 2024), GPT-4o-mini (OpenAI, [n. d.]), and DeepSeek-Coder-V2-Lite-Instruct (abbr. DSCoder) (DeepSeek-AI, 2024). They include both general-purpose LLMs and coding LLMs, and both open-sourced and commercial LLMs from different families. The detailed information of these LLMs is listed in Table 1.
@@
Qwen Open Source 32.8B Mar 2024 (Knowledge Cutoff Information of Qwen32B, [n. d.])
Phi-4 Microsoft Open Source 14.7B May 2024 (Knowledge Cutoff Information of Phi-4, [n. d.])
GPT-4o-mini OpenAI Commercial (Unknown) Oct 2023 (Knowledge Cutoff Information of GPT-4o-mini, [n. d.])
@@
Qwen Open Source 7.62B Mar 2024 (Knowledge Cutoff Information of Qwen7B, [n. d.])
@@
DeepSeek Open Source 15.7B Nov 2023 (Knowledge Cutoff Information of DSCoder, [n. d.])
*: We run Qwen2.5-Coder-32B-Instruct-GPTQ-Int4 on with NVIDIA RTX4090 GPU. The quantized GPTQ-Int4 version is found to show comparable coding ability as the original bf16 version while taking less GPU memory (Wu et al., 2025).
LLM Configuration: We run the four open-source LLMs on our server and access the commercial GPT-4o-mini via OpenAI API. To balance the diversity and reliability, we adopt the temperature of 0.2 following the practice in code generation studies (Jain et al., 2025; Wu et al., 2025; Xu et al., 2024). The maximum output tokens are set to be 3000 (i.e., 1.5x maximum input code length) to accommodate normal code translation outputs. The other configurations (e.g., top_p and penalty) are kept as default values. To mitigate randomness, we repeat the experiments three times and report the average results.
Experimental Environment: We run experiments on our server with Ubuntu 22.04 OS. The server is equipped with NVIDIA RTX4090 GPUs to deploy the open-source LLMs, as well as an AMD Ryzen Threadripper PRO 3995WX 64-Core CPU to support the parallel evaluation of massive generated translation results.
@@ 4. Results and Analysis
4.1. RQ1: Overall Effectiveness of Pseudocode-based Transitive Code Translation
In RQ1, we investigate the overall effectiveness of pseudocode-based code translation. Specifically, we explore (1) how pseudocode-based code translation can help enhance the performance of the current widely-adopted direct translation strategy; and (2) whether pseudocode generated by the translator LLM can effectively represent the code intent or logic of the original program under translation and promote accurate code translation results. We investigate these by comparing the performance of five translation strategies introduced in Section 3.3. We take the widely adopted direct translation strategy ($\mathcal{D}$) as the baseline and analyze the performance improvement of the other four strategies. Table 2 shows the pass@10 rate of each LLM using different strategies.
Table 2. Pass@10 Rate of Code Translation with Different Translation Strategies
@@ 4.1. RQ1: Overall Effectiveness of Pseudocode-based Transitive Code Translation
$\mathcal{PC}$  (viaPseudocode w/ program as context) 0.9648 (+1.77%) 0.9001 (+2.61%) 0.7553 (+7.00%)
$\mathcal{D}\&\mathcal{P}$  (Hybrid [DT & viaPseudo.]) 0.9773 (+3.09%) 0.9403 (+7.19%) 0.7948 (+12.59%)
$\mathcal{D}\&\mathcal{PC}$  (Hybrid [DT & viaPseudo. w/ prog.]) 0.9773 (+3.09%) 0.9327 (+6.33%) 0.7988 (+13.16%)
@@
$\mathcal{PC}$  (viaPseudocode w/ program as context) 0.9254 (-0.24%) 0.8347 (+0.41%) 0.6463 (+2.90%)
$\mathcal{D}\&\mathcal{P}$  (Hybrid [DT & viaPseudo.]) 0.9646 (+3.99%) 0.8947 (+7.63%) 0.7083 (+12.77%)
$\mathcal{D}\&\mathcal{PC}$  (Hybrid [DT & viaPseudo. w/ prog.]) 0.9610 (+3.60%) 0.8903 (+7.10%) 0.7038 (+12.05%)
@@
$\mathcal{PC}$  (viaPseudocode w/ program as context) 0.9239 (-0.43%) 0.8635 (+0.97%) 0.6837 (+2.55%)
$\mathcal{D}\&\mathcal{P}$  (Hybrid [DT & viaPseudo.]) 0.9658 (+4.08%) 0.9054 (+5.87%) 0.7404 (+11.05%)
$\mathcal{D}\&\mathcal{PC}$  (Hybrid [DT & viaPseudo. w/ prog.]) 0.9621 (+3.69%) 0.9144 (+6.92%) 0.7617 (+14.25%)
@@
$\mathcal{PC}$  (viaPseudocode w/ program as context) 0.8989 (-0.93%) 0.7929 (+0.01%) 0.5480 (+1.48%)
$\mathcal{D}\&\mathcal{P}$  (Hybrid [DT & viaPseudo.]) 0.9336 (+2.90%) 0.8395 (+5.89%) 0.5849 (+8.31%)
$\mathcal{D}\&\mathcal{PC}$  (Hybrid [DT & viaPseudo. w/ prog.]) 0.9420 (+3.82%) 0.8568 (+8.07%) 0.6154 (+13.96%)
@@
$\mathcal{PC}$  (viaPseudocode w/ program as context) 0.9218 (+2.83%) 0.7875 (-0.01%) 0.5530 (-1.00%)
$\mathcal{D}\&\mathcal{P}$  (Hybrid [DT & viaPseudo.]) 0.9482 (+5.78%) 0.8569 (+8.80%) 0.6442 (+15.32%)
$\mathcal{D}\&\mathcal{PC}$  (Hybrid [DT & viaPseudo. w/ prog.]) 0.9537 (+6.39%) 0.8581 (+8.95%) 0.6459 (+15.63%)
@@
$\mathcal{PC}$  (viaPseudocode w/ program as context) 0.9270 (+0.60%) 0.8357 (+0.83%) 0.6373 (+2.81%)
$\mathcal{D}\&\mathcal{P}$  (Hybrid [DT & viaPseudo.]) 0.9579 (+3.96%) 0.8874 (+7.06%) 0.6945 (+12.04%)
$\mathcal{D}\&\mathcal{PC}$  (Hybrid [DT & viaPseudo. w/ prog.]) 0.9592 (+4.10%) 0.8905 (+7.44%) 0.7051 (+13.75%)
@@
Effectiveness in Enhancing Direct Translation. The results show that the two hybrid strategies (i.e., $\mathcal{D}\&\mathcal{P}$ and $\mathcal{D}\&\mathcal{PC}$) that combine results from pseudocode-based and direct translation effectively improved the code translation accuracy achieved by the baseline ($\mathcal{D}$). Specifically, $\mathcal{D}\&\mathcal{P}$ improved the pass@10 rate by 3.96%, 7.06%, and 12.04% on average on three difficulty levels, respectively. $\mathcal{D}\&\mathcal{PC}$ led to more significant improvements of 4.10%, 7.44%, and 13.75% on three levels on average, respectively. The results demonstrate that pseudocode-based code translation enables LLMs to correctly translate many programs that cannot be correctly handled by direct translation alone. Thus, given the same number of attempts, it is beneficial to adopt both direct translation and pseudocode-based translation instead of consistently giving chances to either strategy. The results indicate that pseudocode-based code translation can effectively complement the widely adopted direct translation to enhance code translation accuracy.
@@
Meanwhile, we observed that the two pure pseudocode-based translation strategies (i.e., $\mathcal{P}$ and $\mathcal{PC}$) only produce comparable or even worse performance than direct translation ($\mathcal{D}$). Specifically, $\mathcal{PC}$ increased pass@10 of $\mathcal{D}$ by only 0.60%, 0.83%, and 2.81% on three levels on average; $\mathcal{P}$ even led to drops in pass@10 rate by 6.30%, 10.81%, and 16.26%. The results demonstrate that when conducting semantic translation via pseudocode, LLMs also failed to translate a few programs that could be correctly translated by direct translation, indicating that direct translation and pseudocode-based translation are complementary at the current stage. We discuss potential causes of this underperformance via case studies in Section 4.5; we also explore the potential of pure pseudocode-based code translation based on higher-quality pseudocode in RQ4 (Section 4.4).
@@
Difficulty-wise. We also notice that the improvement from hybrid strategies is more significant for harder tasks. Specifically, the improvement in pass@10 rate brought by $\mathcal{D}\&\mathcal{P}$ and $\mathcal{D}\&\mathcal{PC}$ over $\mathcal{D}$ increases by 3.96% and 4.10% on easy-level tasks, while 12.04% and 13.75% on hard-level tasks on average, respectively. The results indicate that pseudocode-based code translation is notably helpful to complement direct translation on relatively complicated programs.
@@
Message 1: Pseudocode-based code translation can effectively complement the widely adopted direct translation strategy to translate programs that cannot be correctly handled by direct translation alone, with a more significant improvement for relatively complicated programs. We recommend practitioners adopt hybrid strategies that combine the strengths of direct translation and pseudocode-based translation to enhance code translation accuracy, by collecting translated programs from both approaches and identifying the final result based on tests.

### v3

@@ 3.1. Research Questions
This RQ aims to show a general picture of the effectiveness of emulating semantic translation for code via pseudocode. We investigate whether pseudocode-based translation effectively complements direct translation and whether it can fully replace direct translation, aiming to shed light on the effective strategy to leverage translation approaches.
RQ2: How helpful is pseudocode-based code translation for different pairs of source and target programming languages?
@@
RQ3: How effective is pseudocode as an intermediary compared to a specific programming language?
Macedo et al. (2025) demonstrates that a specific PL with close syntactic and semantic similarities to the source and target PLs also benefits transitive code translation. For example, translating Python code to Java via a Rust implementation as a bridge is found to complement direct Python-to-Java translation. In this RQ, we aim to understand whether pseudocode, which interprets mainly code intent and logic without PL-specific details, can serve as a more effective intermediary than a specific PL in transitive code translation.
@@
LLMs may not accurately interpret the intent and logic of original programs, which may result in low-quality pseudocode, conversely hindering LLMs from generating correctly translated programs. This RQ investigates the code translation performance based on high-quality pseudocode, aiming to reveal the potential of pseudocode-based code translation and identify the bottlenecks of studied LLMs to harness the potential effectiveness of pseudocode-based code translation.
@@ 3.2.1. Programming Languages
In this study, we evaluate code translation performance across six widely studied PLs, i.e., C++, Python, Java, JavaScript, Go, and Rust. The choice of these PLs follows prior work, where translation among the first four popular PLs (TIOBE Index, ) has been extensively explored (Rozière et al., 2022; Jiao et al., 2023; Yang et al., 2024), while recent studies have highlighted the emerging Go and Rust (Pan et al., 2024; Macedo et al., 2025).
@@ 3.2.2. Code Translation Tasks
We prompt LLMs to translate the solution programs of 323 LeetCode problems in LiveCodeBench-v5 (Jain et al., 2025). We choose these programs because they involve common implementations widely used in practical development and form a representative benchmark to evaluate the coding ability of LLMs (Hui et al., 2024; Team, 2024). Actually, solution programs of LeetCode problems have been widely adopted as subjects in code translation studies, as they provide easy-to-collect multilingual parallel programs (Wang et al., 2023a; Ahmad et al., 2023). Besides, LiveCodeBench offers carefully validated and representative tests, which facilitate the validation of translation results. Additionally, the LiveCodeBench programming problems are released after May 2023 (Jain et al., 2025), allowing the evaluation to be less subject to the data contamination issue compared to the conventional code translation benchmarks like CodeNet (Puri et al., 2021), TransCoder (Rozière et al., 2020), and AVATAR (Ahmad et al., 2023) established in 2020–2021.
Table 1. AST Node Counts and API Invocation Counts of Programs for Different Difficulty-Level Problems
Difficulty Average Count of AST Nodes Average Count of Invocation Nodes
C++ Go Java JS Python Rust C++ Go Java JS Python Rust
Easy
167.91
142.81
159.33
138.50
111.22
173.58
2.56
2.29
2.94
1.97
3.32
5.16
Medium
280.25
263.34
267.83
266.81
188.18
298.54
5.43
4.95
5.46
4.42
5.64
8.69
Hard
592.81
529.85
519.15
546.64
339.70
584.04
10.55
8.85
7.43
11.23
8.77
13.68
To support the investigation of code translation among diverse PLs, we follow existing practices (Jain et al., 2025; Wu et al., 2025) to collect user-written solution programs that can pass all the tests on LiveCodeBench, for all six PLs. These solution programs are used as the original programs to translate. Among all 381 LeetCode problems from LiveCodeBench-v5, we excluded 16 problems with incorrect tests identified by Wu et al. (2025). We also excluded another 42 problems lacking solutions in at least one of the six PLs on LeetCode. This results in 116 easy problems, 160 medium problems, and 47 hard problems. The solution programs of harder problems typically include more complex solving logic and implementations (Jain et al., 2025). As shown in Table 1, the average abstract syntax tree (AST) node count and API invocation count [^1] The statistics are based on counts of all AST nodes and external function calls identified by tree-sitter (Developers, ), respectively. of solution programs for medium/hard-level problems are 1.7x–1.9x / 3.1x–3.9x and 1.7x–2.2x / 2.5x–5.7x as those of easy-level problems across six PLs, suggesting that the harder tasks present higher translation complexity (Ou et al., 2024). Note that, for each problem, LiveCodeBench does not provide corresponding solution programs and only supports the evaluation of Python solutions. We collect the solution programs in six PLs and extend the execution-based evaluation scaffold to all six studied PLs.
@@ 3.3. Experimental Strategies for Code Translation
(a) Prompt for Direct Translation
(b) Prompt for Generating Code from Pseudocode
(c) Prompt for Generating Code from Pseudocode with Original Program as Reference Implementation
@@
    Direct Translation (abbr. $\mathcal{D}$): This strategy prompts LLMs to translate original programs in the source PL to the target PL in one step. The strategy is widely used by existing LLM-based code translation studies (Yang et al., 2024; Pan et al., 2024). We follow these studies to design the prompt as shown in Figure 3(a).
@@
    Transitive Translation via Pseudocode (abbr. $\mathcal{P}$): This strategy prompts LLMs to emulate semantic translation using two transitive steps explicitly: (i) generate a piece of pseudocode to describe the PL-agnostic code intent and logic for the original program; and (ii) implement the pseudocode into a program in the target PL following the prompt in Figure 3(b). As introduced in Section 2.2, this strategy is inspired by human practices of translation and the findings that pseudocode can encode the PL-agnostic code intent and logic of a program (Wu et al., 2025), serving as an appropriate semantic representation for emulating semantic translation. In our study, pseudocode and translated programs are generated by the same LLM (referred to as “translator LLM”). Our prompts are designed following practices of Wu et al. (2025). In particular, we reuse their prompt for pseudocode generation, which guides LLMs to generate pseudocode following: 1) using indents to represent control structures; 2) avoiding usages and type information specific to PLs; 3) omitting implementation of common algorithms, data structures, and operations; and 4) naming functions using the template “function [FUNC_NAME] ([PARAM_LIST])” (e.g., “function max(a, b)”). These conventions align with the general expectations of pseudocode style in prior studies (Kulal et al., 2019; Kulkarni et al., ; Acharjee et al., 2022). Given that there is no unified format and grammar for pseudocode, we allow LLMs to generate pseudocode in their most familiar style based on these conventions. This prompt is available in our artifact (Artifact of this paper, ).
@@
    Transitive Translation via Pseudocode with Original Program as Context (abbr. $\mathcal{PC}$): This strategy extends the above pseudocode-based strategy to include implementation details (i.e., original program) as the context to guide code translation. Specifically, considering that pseudocode abstracts implementation details of the original program, it may miss some helpful information for LLMs to implement the original program’s intent and logic in the target PL. Thus, we include the original program in the source PL as an implementation example of the pseudocode for LLMs’ reference when generating the translated program by emulating semantic translation, using the prompt shown in Figure 3(c).
@@
We also examine the effectiveness of combining direct translation and pseudocode-based translation. Specifically, we design two straightforward hybrid strategies by evenly mixing direct translation results and pseudocode-based translation results. We keep the same total attempts of ten as the above three single-approach strategies, thus combining five translated programs generated by each approach. The two hybrid strategies are defined as follows:
@@ 3.4. Experimental Setups
Metric: We report the pass@10 rate of the translated programs (i.e., programs in the target PL) on the test cases provided by LiveCodeBench to reflect the computational accuracy of the translation results. The metric measures the ratio of tasks on which the translator LLM successfully generates a correct translation in at least one out of ten attempts. Compared with text-based metrics like BLEU and Exact Match rate, the execution-based pass@10 metric can measure the semantic equivalence of the translation results based on the tests provided by LiveCodeBench. The metric is widely adopted in code generation (Jain et al., 2025; Wu et al., 2025) and translation studies (Macedo et al., 2025). Note that we use pass@10 to investigate the potential of the studied translation strategies and LLMs, considering the higher translation accuracy achieved with ten attempts than one single attempt demonstrated by existing study (Macedo et al., 2025) and the practical feasibility of filtering translation results by referring to the original program’s behavior as we discuss in Section 2.1. In addition, we discuss pass@{2,4,6,8} in RQ1.
Studied LLMs: In this study, we investigate five popular LLMs deployable on our machine. They are Qwen2.5-Coder-32B-Instruct (abbr. Qwen32B) (Hui et al., 2024), Qwen2.5-Coder-7B-Instruct (abbr. Qwen7B) (Hui et al., 2024), Phi-4 (Abdin et al., 2024), GPT-4o-mini (OpenAI, ), and DeepSeek-Coder-V2-Lite-Instruct (abbr. DSCoder) (DeepSeek-AI, 2024). They include both general-purpose LLMs and coding LLMs, and both open-sourced and commercial LLMs from different families. The detailed information of these LLMs is listed in Table 2.
@@
Qwen Open Source 32.8B Mar 2024 (Knowledge Cutoff Information of Qwen32B, )
Phi-4 Microsoft Open Source 14.7B May 2024 (Knowledge Cutoff Information of Phi-4, )
GPT-4o-mini OpenAI Commercial (Unknown) Oct 2023 (Knowledge Cutoff Information of GPT-4o-mini, )
@@
Qwen Open Source 7.62B Mar 2024 (Knowledge Cutoff Information of Qwen7B, )
@@
DeepSeek Open Source 15.7B Nov 2023 (Knowledge Cutoff Information of DSCoder, )
*: We run Qwen2.5-Coder-32B-Instruct-GPTQ-Int4 with NVIDIA RTX4090 GPU. The quantized GPTQ-Int4 version is found to show comparable coding ability as the original bf16 version while taking less GPU memory (Wu et al., 2025).
LLM Configuration: We run the four open-source LLMs on our machine and access the commercial GPT-4o-mini via OpenAI API. To balance the diversity and reliability, we adopt the temperature of 0.2 following the practice in LLM-based coding studies (Jain et al., 2025; Wu et al., 2025; Xu et al., 2024). The maximum output tokens are set to be 3000 (representing 1.5x maximum input code length) to accommodate normal translation outputs. The other configurations (e.g., top_p and penalty) are kept as default values. To mitigate randomness, we repeat the experiments three times and report the average results.
Experimental Environment: We run experiments on a machine with Ubuntu 22.04 OS. The server is equipped with NVIDIA RTX4090 GPUs to deploy the open-source LLMs, as well as an AMD Ryzen Threadripper PRO 3995WX 64-Core CPU to support the parallel evaluation of plenty of generated translation results.
@@ 4. Results and Analysis
4.1. RQ1: Overall Effectiveness of Pseudocode-based Code Translation
In RQ1, we investigate the overall effectiveness of pseudocode-based code translation. Specifically, we explore (1) whether pseudocode-based code translation can advance the accuracy of the current widely-adopted direct translation approach, by either complementarily enhancing or completely replacing direct translation; and (2) whether pseudocode generated by the translator LLM can effectively represent the code intent or logic of the original program under translation and promote accurate code translation results. We investigate these by comparing the performance of five translation strategies introduced in Section 3.3. We take the widely adopted direct translation strategy ($\mathcal{D}$) as the baseline and analyze the performance improvement of the other four strategies. Table 3 shows the pass@10 rate of each LLM using different strategies.
Table 3. Pass@10 Rate of Code Translation Using Different Translation Strategies
@@ 4.1. RQ1: Overall Effectiveness of Pseudocode-based Code Translation
$\mathcal{PC}$  (viaPseudocode w/ original program as context) 0.9648 (+1.77%) 0.9001 (+2.61%) 0.7553 (+7.00%)
$\mathcal{D}\&\mathcal{P}$  (Hybrid [DT & viaPseudocode]) 0.9773 (+3.09%) 0.9403 (+7.19%) 0.7948 (+12.59%)
$\mathcal{D}\&\mathcal{PC}$  (Hybrid [DT & viaPseudocode w/ context]) 0.9773 (+3.09%) 0.9327 (+6.33%) 0.7988 (+13.16%)
@@
$\mathcal{PC}$  (viaPseudocode w/ original program as context) 0.9254 (-0.24%) 0.8347 (+0.41%) 0.6463 (+2.90%)
$\mathcal{D}\&\mathcal{P}$  (Hybrid [DT & viaPseudocode]) 0.9646 (+3.99%) 0.8947 (+7.63%) 0.7083 (+12.77%)
$\mathcal{D}\&\mathcal{PC}$  (Hybrid [DT & viaPseudocode w/ context]) 0.9610 (+3.60%) 0.8903 (+7.10%) 0.7040 (+12.08%)
@@
$\mathcal{PC}$  (viaPseudocode w/ original program as context) 0.9239 (-0.43%) 0.8635 (+0.97%) 0.6837 (+2.55%)
$\mathcal{D}\&\mathcal{P}$  (Hybrid [DT & viaPseudocode]) 0.9658 (+4.08%) 0.9054 (+5.87%) 0.7404 (+11.05%)
$\mathcal{D}\&\mathcal{PC}$  (Hybrid [DT & viaPseudocode w/ context]) 0.9621 (+3.69%) 0.9144 (+6.92%) 0.7617 (+14.25%)
@@
$\mathcal{PC}$  (viaPseudocode w/ original program as context) 0.8989 (-0.93%) 0.7929 (+0.01%) 0.5480 (+1.48%)
$\mathcal{D}\&\mathcal{P}$  (Hybrid [DT & viaPseudocode]) 0.9336 (+2.90%) 0.8395 (+5.89%) 0.5849 (+8.31%)
$\mathcal{D}\&\mathcal{PC}$  (Hybrid [DT & viaPseudocode w/ context]) 0.9420 (+3.82%) 0.8568 (+8.07%) 0.6154 (+13.96%)
@@
$\mathcal{PC}$  (viaPseudocode w/ original program as context) 0.9218 (+2.83%) 0.7875 (-0.01%) 0.5530 (-1.00%)
$\mathcal{D}\&\mathcal{P}$  (Hybrid [DT & viaPseudocode]) 0.9482 (+5.78%) 0.8569 (+8.80%) 0.6442 (+15.32%)
$\mathcal{D}\&\mathcal{PC}$  (Hybrid [DT & viaPseudocode w/ context]) 0.9537 (+6.39%) 0.8581 (+8.95%) 0.6459 (+15.63%)
@@
$\mathcal{PC}$  (viaPseudocode w/ original program as context) 0.9270 (+0.60%) 0.8357 (+0.83%) 0.6373 (+2.81%)
$\mathcal{D}\&\mathcal{P}$  (Hybrid [DT & viaPseudocode]) 0.9579 (+3.96%) 0.8874 (+7.06%) 0.6945 (+12.04%)
$\mathcal{D}\&\mathcal{PC}$  (Hybrid [DT & viaPseudocode w/ context]) 0.9592 (+4.10%) 0.8905 (+7.44%) 0.7052 (+13.76%)
@@
Effectiveness in Enhancing Direct Translation. The results show that the two hybrid strategies (i.e., $\mathcal{D}\&\mathcal{P}$ and $\mathcal{D}\&\mathcal{PC}$) that combine translation outcomes from pseudocode-based and direct translation effectively improved the code translation accuracy achieved by the baseline ($\mathcal{D}$). Specifically, $\mathcal{D}\&\mathcal{P}$ improved the pass@10 rate by 3.96%, 7.06%, and 12.04% on average on three difficulty levels, respectively. $\mathcal{D}\&\mathcal{PC}$ led to more significant improvements of 4.10%, 7.44%, and 13.76% on three levels on average, respectively. The results demonstrate that pseudocode-based code translation enables LLMs to correctly translate many programs that cannot be correctly handled by direct translation alone. Thus, given the same number of attempts, it is beneficial to adopt both direct translation and pseudocode-based translation instead of consistently giving chances to either strategy. The results indicate that pseudocode-based code translation can effectively complement the widely adopted direct translation to enhance code translation accuracy.
@@
Meanwhile, we observed that the two pure pseudocode-based translation strategies (i.e., $\mathcal{P}$ and $\mathcal{PC}$) produced only comparable or even worse performance than direct translation ($\mathcal{D}$). Specifically, $\mathcal{PC}$ increased pass@10 of $\mathcal{D}$ by only 0.60%, 0.83%, and 2.81% on three levels on average; $\mathcal{P}$ even led to drops in pass@10 rate by 6.30%, 10.81%, and 16.26%. The results demonstrate that when conducting semantic translation via pseudocode, LLMs also failed to translate a few programs that could be correctly translated by direct translation. This indicates that direct translation and pseudocode-based translation are complementary at the current stage. We discuss how this underperformance can be attributed to incorrect, incomplete, and ambiguous pseudocode via case studies in Section 4.5, and we also explore the potential of pure pseudocode-based code translation based on higher-quality pseudocode in RQ4 (Section 4.4).
@@
Difficulty-wise. We also notice that the improvement of hybrid strategies is more significant on harder tasks. Specifically, the improvement in pass@10 rate brought by $\mathcal{D}\&\mathcal{P}$ and $\mathcal{D}\&\mathcal{PC}$ over $\mathcal{D}$ increases by 3.96% and 4.10% on easy-level tasks, while 12.04% and 13.76% on hard-level tasks on average, respectively. As introduced in Section 3.2, the original programs of harder LeetCode problems typically include more complex solving logic and implementations, involving more AST nodes and more external API invocations to translate. The results indicate that pseudocode-based code translation is notably helpful in complementing direct translation for relatively complicated programs. Our case studies in Section 4.5.1 will further discuss the advantages of pseudocode-based translation, which also align with this observation.
Figure 4. Pass Rates of Direct Translation $\mathcal{D}$ and Hybrid Strategy $\mathcal{D}\&\mathcal{PC}$ with Different Numbers of Attempts
[…]
@@
Message 1: Pseudocode-based code translation can effectively complement the widely adopted direct translation approach to translate programs that cannot be correctly handled by direct translation alone, with a more significant improvement for more complicated programs. We recommend that practitioners adopt hybrid strategies that combine the strengths of direct translation and pseudocode-based translation to enhance translation accuracy.